y"−3y' + 2y = 3e−x−10cos3x, y(0) = 1, y'(0) = 2.
1
Expert's answer
2020-04-09T14:53:21-0400

The general solution of the second order nonhomogeneous linear equation can be expressed in the form


"y=y_c+Y"

where "Y" is any specific function that satisfies the nonhomogeneous equation, and "y_c=C_1y_1+C_2y_2" is a general solution of the corresponding homogeneous equation


"y''-3y'+2y=0"

The Characteristic Equation


"r^2-3r+2=0""(r-1)(r-2)=0"

"r_1=1, r_2=2"

The general solution of the corresponding homogeneous equation is


"y_c=C_1e^{x}+C_2e^{2x}"

Use Method of Undetermined Coefficients


"Y=Ae^{-x}+B\cos(3x)+C\sin(3x)"


"Y'=-Ae^{-x}-3B\sin(3x)+3C\cos(3x)"

"Y''=Ae^{-x}-9B\cos(3x)-9C\sin(3x)"

"Ae^{-x}-9B\cos(3x)-9C\sin(3x)+"

"+3Ae^{-x}+9B\sin(3x)-9C\cos(3x)+"

"+2Ae^{-x}+2B\cos(3x)+2C\sin(3x)=3e^{-x}-10\cos(3x)"

"6A=3=>A=\dfrac{1}{2}"

"-7B-9C=-10"

"9B-7C=0"

"B=\dfrac{7}{13} , C=\dfrac{9}{13}"


"Y=\dfrac{1}{2}e^{-x}+\dfrac{7}{13}\cos(3x)+\dfrac{9}{13}\sin(3x)"


"y=C_1e^{x}+C_2e^{2x}+\dfrac{1}{2}e^{-x}+\dfrac{7}{13}\cos(3x)+\dfrac{9}{13}\sin(3x)"

"y(0)=1:"

"1=C_1+C_2+\dfrac{1}{2}+\dfrac{7}{13}"

"y'(0)=2:"

"y'=C_1e^{x}+2C_2e^{2x}-\dfrac{1}{2}e^{-x}-\dfrac{21}{13}\sin(3x)+\dfrac{27}{13}\cos(3x)"

"2=C_1+2C_2-\dfrac{1}{2}+\dfrac{27}{13}"

"1=C_2-1+\dfrac{20}{13}"

"C_2=\dfrac{6}{13}"


"C_1=-\dfrac{1}{2}"

"y_p=-\dfrac{1}{2}e^{x}+\dfrac{6}{13}e^{2x}+\dfrac{1}{2}e^{-x}+\dfrac{7}{13}\cos(3x)+\dfrac{9}{13}\sin(3x)"